OpenStudy (anonymous):

The first four terms of a sequence are shown below: 6, 2, -2, -6 Which of the following functions best defines this sequence? f(1) = 6, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 2; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) + 2; for n ≥ 1

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

forget the answer choices for a second

jimthompson5910 (jim_thompson5910):

focus on 6, 2, -2, -6 what rule is being applied here?

OpenStudy (anonymous):

*-3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so you're saying "multiply by -3" ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

6 times -3 = -18 though

jimthompson5910 (jim_thompson5910):

I start with 6 and I want to get to 2 what operation can I do to get there?

OpenStudy (anonymous):

minus

OpenStudy (anonymous):

take away @jim_thompson5910

jimthompson5910 (jim_thompson5910):

6 minus _______ = 2 what goes in the blank

OpenStudy (anonymous):

4!

jimthompson5910 (jim_thompson5910):

so we're subtracting 4 from each term to get the next term

jimthompson5910 (jim_thompson5910):

we have 2, we want to get to -2 2 - 4 = -2 then we subtract 4 again -2 - 4 = -6 etc etc

jimthompson5910 (jim_thompson5910):

so the rule is f(1) = 6, f(n + 1) = f(n) - 4; for n ≥ 1 which basically says "the first term is 6. To get any new terms, subtract 4 from the last term" f(n+1) is the next term after term f(n)

OpenStudy (anonymous):

wow that sounds so confusing

jimthompson5910 (jim_thompson5910):

it's not too bad once you get the hang of it

jimthompson5910 (jim_thompson5910):

it's like saying f(3) = f(2) - 4

OpenStudy (anonymous):

yea now i see that -4 and that simplifies it hahah

jimthompson5910 (jim_thompson5910):

I'm glad it does

OpenStudy (anonymous):

thanks lol do you still have time? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes I do